# Helfrich cylinder Alexander Meiners, Hannes Uecker

### v1

@@ 5.2 $c_{0}=-1$
We also ran some numerical flows from perturbations of A in Fig.12, and similarly at larger $L$, with results as in Fig.11, i.e., convergence to the single pearl solutions. However, these DNS are very expensive, partly due to the larger domain and hence larger scale discretization, but more importantly due to the very slow convergence of the flows.
[…]
Figure 12: $c_{0}=0.48$, $L=30$. The $m=4$ (period $\ell=7.5$) pearling A now looses stability at BP1 much closer to its primary bifurcation, to a single “localized” pearling branch, sample B. Further branches: C: $m=2$ ($\ell=15$) primary pearling; E: $m=3$ ($\ell=10)$ primary pearling; D: 2ndary pearling from 2nd BP on $m=1$ branch. F: primary $(m,n)=(1,1)$ coiling.
Appendix A Appendix
A.1 Numerical continuation via pde2path
@@
The stability analysis here is done in a slightly different setting as in the legacy version of pde2path. Usually the stability is checkt for the given active parameter set p.nc.ilam.m, in our case $(\lambda_{1},\lambda_{2},\text{PC})$, where PC stands for the Lagrange multipliers of the phase conditions, where the primary $\lambda_{1}$ is assumed to be fixed. However in the dynamical problem, the Lagrange multipliers are in general time dependent. The disinherit, stability information is for fixed $A$ and $V$. Hence after a successful continuation step, a new Jacobian is build according to the p.nc.ilam$=(V,\lambda_{1},\lambda_{2},\text{PC})$.
@@
During the continuation of bifurcating branches, it is not necessary that the boundary of $X$ stays in a plane, similar to the cylinder. This however is allowed and makes no problem in the continuation. But for mesh refinement along the branch the legendary box2per setting looks for periodic boundary nodes in a plane. For that reason we only allow an $(y,z)$ displacement at the boundary, by setting $N_{x}=0$. Note that the computations for $H(X)$ this is not done for accurate computations also at the boundary.
A.1.1 Mesh handling
The initial cylinder in discretized by a standard mesh from pde2path, mapped via a local map on the cylinder. The number of nodes in $x$ and $y$ direction is chosen that the quotient
$q=\frac{e_{M}}{R},$ (73)
with $e_{M}$ the maximal edge length and $R=A/C$ the ratio of area and circumcenter, is minimal. The triangulation is oriented inwards, matching with the above convention.
On the bifurcation branches some triangles may distort, due to the change of X. We establish an mesh adaption consisting of refinement and coarsening of X and tri from the prior step wrt. the periodic boundary nodes. As a control function we use the are of a triangle and refine the area is above a threshold. This is typically 1.3 times the size of a triangle on the original cylinder.The refinement algorithm Trefinelong.m is based on the RGB refinement in [FS21], but force that the longest side has to be refined, see LABEL:reffig(b) , creating a mesh with valences 4 and 8. However this increases (73) in a regular sized mesh, while for acute triangles the refinement produces two obtuse ones. A retriangulation retrigX.m swaps edges so we get only valence 5 or 6 and lowers (73) again, see LABEL:reffig(c) , except at the boundary, see [PS04]. However this strategy has proven to be the most efficient way, in order to avoid long cascades of triangles from the RGB refinement. For the conservation of the periodic boundary, we enforce the same refinement at both boundary triangles. In some cases one boundary side is refined while the local structure do not allow the same refinement at a the other boundary. In that case we collapse the excess node along the side of its nearest boundary neighbor.
For the coarsening (73) is the control parameter, activated if $q>8$. The strategy select triangles smaller then a threshold $A_{t}$, (legacy version max(A)/2, but set to max(A)/3 and acute (ratio of circumcenter to area) as coarsened triangles. Then the obtuse neighbor triangles of those are found. The acute triangles are collapsed along the shortest edge. Then edges are flipped until one gets a triangular mesh. This is iterated as long as all triangles are larger than $A_{t}$.

### v2

@@ 5.2 $c_{0}=-1$
We also ran some numerical flows from perturbations of A in Fig.13, and similarly at larger $L$, with results as in Fig.12, i.e., convergence to the single pearling solutions. However, these DNS are very expensive, partly due to the larger domain and hence larger scale discretization, but more importantly due to the very slow convergence of the flows.
5.4 Outlook: $c_{0}=0.75$
Lastly, following Fig.2(e) we consider $c_{0}>1/2$, specifically $c_{0}=0.75$ for which ${\cal C}_{L}$ is never stable if $L\geq 2\pi$ ($k_{1}\leq 1$, say), and branches as before still bifurcate from ${\cal C}_{L}$, but not at loss of stability of ${\cal C}_{L}$. Hence they can also still be described by AEs as before, but they cannot be stable close to bifurcation, and we can only expect stable “patterns” at $O(1)$ amplitude. This is more difficult than the analysis and numerics so far, and hence here we only give an outlook on some interesting (numerical) results we obtain, again for a rather small length $L=10$ of the cylinder, and leave more systematic and detailed studies to later work.
Figure 14 shows the BD of $k_{1}\approx 0.628$ pearling (A, blue) and $k_{2}\approx 1.257$ pearling (C, light blue), which are both unstable at bifurcation as they must, and moreover are unstable throughout, and addtionally an $O(1)$ amplitude pearling branch (P, red), where P2 is obtained from DNS starting near A1 (panel (c)). In summary, we see a number of new effects which strongly suggest further study, but here we restrict to the following comments.
(a) (b)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Selected branches for $c_{0}=0.75$, $L=10$. (a) Energy $E$ over reduced volume $v$, and $\lambda_{1}$ over $\lambda_{2}$. (b) Samples from primary pearling (blue branch in (a)) all unstable. (c) DNS from near $A_{2}$, converging to a stable pearling $P_{2}$ with unequal pearl sizes. (d) Samples from the stable P branch, and sample from (unstable) primary $m=2$ pearling branch (light blue in (a)).
$\bullet$ We stop the continuations shortly after A3, shortly after C, and shortly after P1 as further continuations there require excessive mesh adaptations. We can continue significantly beyond P3, but our main region of interest is near $v=1$. There are bifurcations on the blue branches which break the rotational symmetry, but we restrict to presenting axisymmetric branches.
$\bullet$ The $k_{1}$ pearling branch A shows two folds in $v$, the first at $v_{f,1}\approx 1.045$, the second at $v_{f,2}\approx 0.98$. In particular, between A1 and A2 it goes through $v=1$ again, while keeping the area $\cal A$} constant throughout. This is not a contradiction to the fact that normal variations of ${\cal C}_{L}$ that fix $\cal A$} and $\cal V$} locally do not exist; the passage through $v=1$ here corresponds to a non–small normal variation.
$\bullet$ The folds on the A branch naturally call for fold continuation in $c_{0}$, in particular for decreasing $c_{0}$ as for $c_{0}<1/2$ we do not see folds on the analogous A branches. We believe that they vanish at some $c_{0}>1/2$ as the blue swallowtail shrinks and the folds collide, but this remains to be studied.
$\bullet$ The flow in (c) starts at A1 with a small pertubation in direction of the most unstable eigenvalue; A1 is below the fold, and there is exactly one positive eigenvalue $\mu\approx 0.2$. In any case, the results of flow experiments seem rather robust wrt the choice of the initial conditions (choice of steady state and perturbation), i.e., we always “land” on the P branch, though naturally at different $v$ and $\lambda_{2}$, or analoga of the P branch at different area $\cal A$}. In the particular case shown, the perturbation of A1 yields a relative change of $\cal V$} of order $10^{-3}$, and a relative change of $\cal A$} of order $10^{-4}$. (The initial perturbation linearly conserves $\cal A$} and $\cal V$}, i.e., is tangential to the constraint manifold, but for any finite perturbation we must expect a (nonlinear) change of $\cal A$} and $\cal V$}. However, the change in $\cal A$} is so small that we ignore it and keep the normalization of $E$ in (71) by the original $\cal A$}.)
$\bullet$ The last plots in (c) suggest almost steady states for $t>20$, and we use the solution at $t=20.95$ as an initial guess for Newton’s method for the steady problem. This indeed takes us to the (spectrally) stable steady state P2, and continuing from here to larger/smaller $v$ yields the red branch.
$\bullet$ It is remarkable that the P branch with two pearls of different sizes is stable, that along the branch the pearls change their relative sizes (different from the primary pearling branches, where only the global amplitude changes), and that in fact P2 attracts (perturbations of) A1. However, some finite size effects are involved here, because some stability of the P branch is lost when repeating the experiment from Fig.14 on larger domains, e.g., $L=20$. In this case, initial conditions for numerical flows must be chosen more carefully, namely closer to $v=1$ to reach the analogue of the P branch (4 pearls on $L=20$), and continuation of the P branch then shows a loss of stability (wrt to still axially symmetric solutions) away from $v=1$. We plan to study in particular these effects in a systematic way in future work, including their relation to the transition from A to B in Fig. 13.
$\bullet$ Shape wise (compare P1 and C in (d)) it seems possible that the P branch connects to the C branch in a period–doubling bifurcation, but samples P1 and C are at rather different values of $\lambda_{1,2}$. In our preliminary experiments so far we could not identify to what other branches the P branch connects, and again we leave this for future work.
6 Discussion
Motivated by experimental results as in Fig.1(a,b), we studied the bifurcations from straight cylinders ${\cal C}_{L}$ of lengths $L$ (and wlog radius $r=1$) in the Helfrich model (5). After showing well–posedness we discussed the linearization at ${\cal C}_{L}$ and derived the amplitude equations for the bifurcations of pearling, wrinkling, and coiling and buckling. Subsequently we used numerical continuation and bifurcation methods to extend the local bifurcations to a more global picture.
Our results initially suggest that (5) besides being very interesting mathematically is a possible model for the experiments in Fig.1(a,b). However, when restricting to branches bifurcating at loss of stability of ${\cal C}_{L}$, there also are strong discrepancies. One is that coiling (and buckling, which as far as we know is not clearly documented in experiments) always bifurcate with the maximum wavelength $L$ allowed by the domain. For $c_{0}>1/4$, pearling can occur with a finite wavelength (Figs.1113), but the stability range of finite wavelength pearling shrinks with increasing $L$, probably to zero for $L\to\infty$. This means that the experimental results in Fig.1(a,b) are not described by the primary branches bifurcating at loss of stability of ${\cal C}_{L}$. On the other hand, when further increasing $c_{0}$ beyond $c_{0}{=}1/2$ we find stable finite wavelength $O(1)$ amplitude pearling, but we only gave an outlook on this in §5.4, and this clearly needs to be studied more systematically and in more detail, and for larger $L$.
Mathematically, the bifurcations of localized pearling from finite wave number pearling, with the pertinent BP moving closer to the primary bifurcation for larger domain size $L$, is similar to the bifurcation of localized patterns from periodic patterns in various dissipative systems, e.g., reaction diffusion (RD) systems or Swift–Hohenberg (SH) type equations. See, e.g., [BK07, HK09, BKL+09, Kno15], and [Uec21, Chapters 8 and 9] for further references and numerics using pde2path. In these systems, the localized patterns are usually attributed to subcritical primary bifurcations of periodic patterns, which are hence unstable, but stabilize after a fold at large amplitude. The localized patterns then show some “snaking” behavior with several folds in which the pattern grows, and with alternating stable and unstable segments of this snaking branch. Moreover, in the limit $L\to\infty$, the BP of the snaking branch converges to the primary BP of the periodic patterns.
The data in (76) for BP1 moving closer to $v=1$ for larger $L$ suggest a similar effect, but there are clear differences to the snaking scenario, and many open questions. In our numerics:
  • After the loss of stability of the primary pearling A–branch, it never regains stability; in particular, we cannot find a fold (in $v$, say) for the A–branch.
  • Similarly, we cannot find a fold or other loss of stability of the localized pearling B–branch.
The numerics here are harder for large $L$ than for periodic or localized 1D (or 2D) patterns in, e.g., reaction diffusion systems, and for instance for $L=60$ we only continue the A–branch and the B–branch to $v=1.01$, such that these results are in some sense still local. As mentioned in §5.4, one way to proceed should be fold continuation in $c_{0}$ of the folds that we find on the A branch at larger $c_{0}$.
All our numerics and associated analysis was for finite (and mostly rather small) $L$. For linearizations of (classical) dissipative systems on intervals $(0,L)$ at some steady state, the spectral gap between the critical eigenvalues and the rest of the spectrum shrinks as $L\to\infty$, and over $\mathbb{R}$} the amplitude equations (ODEs) for finite $L$ must be replaced by modulation equations (PDEs) of Ginzburg–Landau type, incorporating the continuous curve of critical eigenvalues. See, e.g., [SU17, Part IV], and the references therein. For the Helfrich cylinders it is not clear how to proceed for $L\to\infty$. For long wave instabilities (buckling, coiling, wrinkling, and pearling except in a small range of positive $c_{0}$ as in Fig.1113), without the volume constraint the critical curve ${\mathbb{R}}\ni k\mapsto\mu(k,n)$ is maximal at wave number $k=0$, and we believe that instead of (5) a better way to incorporate the volume constraint is to consider the $H^{-1}$ gradient flow for (1) with area constraint. This automatically conserves $\cal V$}, and the critical curve is changed to $k\mapsto\tilde{\mu}(k,n)=k^{2}\mu(k,n)$, such that we expect modulation equations with an (additional) conservation law; see, e.g., [MC00, ZS16, Hil20]. This, however, is also beyond the scope of the present paper.
Appendix A Buckling and coiling amplitude system computations
To derive the coefficients $a,b_{1},b_{2}$ for the amplitude system (67) for the mixed mode instability we as usual plug the extended ansatz (66) into (51) and sort wrt powers of $\varepsilon$.
  • $\varepsilon\psi_{11}$:
    $-\frac{1}{2}k^{2}\left(k^{2}-4c_{0}-2\lambda_{2}+2\right)A=0.$
  • $\varepsilon\psi_{1-1}$:
    $-\frac{1}{2}k^{2}\left(k^{2}-4c_{0}-2\lambda_{2}+2\right)B=0$.
  • $\varepsilon^{2}\psi_{00}$:
    $0=\left(-\frac{1}{2}+\frac{k^{4}}{2}+\frac{\left(5-8c_{0}-2\lambda_{2}\right)k^{2}}{2}-\lambda_{2}\right)({|A|}^{2}+{|B|}^{2})-(\lambda_{2}+1/2)A_{00}-\beta_{2}-\alpha_{2}.$
  • $\varepsilon^{2}\psi_{22}$:
    $0=\left(-\frac{7k^{4}}{4}{+}\frac{\left(-8c_{0}{+}2\lambda_{2}{+}19\right)k^{2}}{4}{-}\frac{3\lambda_{2}}{2}{+}\frac{9}{4}\right)A^{2}+\left(-8k^{4}{+}\frac{\left(16\lambda_{2}{+}32c_{0}{-}64\right)k^{2}}{4}{+}3\lambda_{2}{-}\frac{9}{2}\right)A_{22}.$
  • $\varepsilon^{2}\psi_{2-2}$:
    $0=\left(-\frac{7k^{4}}{4}+\frac{\left(-8c_{0}{+}2\lambda_{2}{+}19\right)k^{2}}{4}-\frac{3\lambda_{2}}{2}{+}\frac{9}{4}\right)B^{2}+\left(-8k^{4}{+}\frac{\left(16\lambda_{2}{+}32c_{0}{-}64\right)k^{2}}{4}+3\lambda_{2}{-}\frac{9}{2}\right)B_{22}$.
  • $\varepsilon^{2}\psi_{20}$:
    $0=\left(-\frac{7k^{4}}{2}{+}\frac{\left(2\lambda_{2}+8c_{0}+3\right)k^{2}}{2}{-}\lambda_{2}{-}\frac{1}{2}\right)BA+\left(-8k^{4}+\frac{\left(16c_{0}{+}8\lambda_{2}\right)k^{2}}{2}{-}\lambda_{2}{-}\frac{1}{2}\right)A_{20}$.
  • $\varepsilon^{2}\psi_{02}$:
    $0=4A\left(\frac{k^{4}}{8}+\left(c_{0}-\frac{\lambda_{2}}{4}+\frac{5}{8}\right)k^{2}-\frac{3\lambda_{2}}{4}+\frac{9}{8}\right)\overline{B}+\left(3\lambda_{2}-\frac{9}{2}\right)A_{02}$.
  • $\varepsilon^{3}\psi_{11}$:
    ${\hskip-28.45274pt}\frac{{\rm d}}{{\rm d}T}A=$ $\bigg{[}(-k^{2}\alpha_{2}-4\left(-\frac{5k^{6}}{8}+\left(c_{0}+\frac{3\lambda_{2}}{8}-2\right)k^{4}+\left(-\frac{c_{0}}{2}+\frac{5\lambda_{2}}{8}+\frac{23}{16}\right)k^{2}-\frac{3\lambda_{2}}{4}+\frac{9}{8}\right){|A|}^{2}$
    ${-}8\left({-}\frac{5k^{6}}{8}{+}\left(c_{0}{+}\frac{3\lambda_{2}}{8}\right)k^{4}{+}\left(\frac{3c_{0}}{2}{-}\frac{3\lambda_{2}}{8}{+}\frac{15}{16}\right)k^{2}{-}\frac{3\lambda_{2}}{4}{+}\frac{9}{8}\right){|B|}^{2}{-}4k^{2}\left(\frac{1}{2}c_{0}{-}\frac{3}{8}\right)A_{00}\bigg{]}A$
    ${+}6\left(\left(c_{0}{+}\frac{1}{4}\right)k^{2}{-}\left(\lambda_{2}{-}\frac{3}{2}\right)\right)A_{02}B{-}10\left(-\frac{2k^{4}}{5}{+}\left(c_{0}{+}\frac{\lambda_{2}}{5}{-}\frac{49}{20}\right)k^{2}{+}\frac{3\lambda_{2}}{5}{-}\frac{9}{10}\right)A_{22}\overline{A}$
    $-2k^{2}\left(-2k^{2}+c_{0}+\lambda_{2}-\frac{1}{4}\right)A_{02}\overline{B}.$
  • $\varepsilon^{3}\psi_{1-1}$:
    ${\hskip-28.45274pt}\frac{{\rm d}}{{\rm d}T}B=$ $\bigg{[}{-}k^{2}\alpha_{2}{-}4\left({-}\frac{5k^{6}}{8}{+}\left(c_{0}{+}\frac{3\lambda_{2}}{8}{-}2\right)k^{4}{+}\left(-\frac{c_{0}}{2}{+}\frac{5\lambda_{2}}{8}{+}\frac{23}{16}\right)k^{2}{-}\frac{3\lambda_{2}}{4}{+}\frac{9}{8}\right){|B|}^{2}$
    ${-}8\Big{(}{-}\frac{5k^{6}}{8}{+}\left(c_{0}{+}\frac{3\lambda_{2}}{8}\right)k^{4}{+}\left(\frac{3c_{0}}{2}{-}\frac{3\lambda_{2}}{8}{+}\frac{15}{16}\right)k^{2}{-}\frac{3\lambda_{2}}{4}{+}\frac{9}{8}\Big{)}{|A|}^{2}{-}2k^{2}\left(c_{0}{-}\frac{3}{4}\right)A_{00}\bigg{]}B$
    $+6\left(\left(c_{0}{+}\frac{1}{4}\right)k^{2}{-}\lambda_{2}{+}\frac{3}{2}\right)A\overline{A_{02}}{-}10\bigg{(}{-}\frac{2k^{4}}{5}{+}\left(c_{0}{+}\frac{\lambda_{2}}{5}{-}\frac{49}{20}\right)k^{2}+\frac{3\lambda_{2}}{5}-\frac{9}{10}\bigg{)}B_{22}\overline{B}$
    ${-}2k^{2}\left({-}2k^{2}{+}c_{0}{+}\lambda_{2}{-}\frac{1}{4}\right)A_{20}\overline{A}.$
From the area constraint we get, at $\varepsilon^{2}\psi_{00}$, $0=\left(k^{2}+1\right)(|A|^{2}+|B|^{2})+A_{00}$, and using $\lambda_{2}^{(1,1)}=\frac{1}{2}k^{2}-2c_{0}+1$ we solve the algebraic equations for the $\varepsilon^{2}$ modes:
$A_{00}=-\frac{\left(4k^{2}c_{0}-2k^{2}-4c_{0}+3\right)}{k^{2}-4c_{0}+3}(|A|^{2}+|B|^{2})-\frac{2\alpha_{2}+2\beta_{2}}{k^{2}-4c_{0}+3},$
$A_{22}=-\frac{\left(2k^{4}+4k^{2}c_{0}-6k^{2}-4c_{0}-1\right)}{2\left(4k^{4}+7k^{2}+4c_{0}+1\right)}A^{2},\quad B_{22}=-\frac{\left(2k^{4}+4k^{2}c_{0}-6k^{2}-4c_{0}-1\right)}{2\left(4k^{4}+7k^{2}+4c_{0}+1\right)}B^{2},$
$A_{20}=-\frac{\left(6k^{4}-4k^{2}c_{0}-4k^{2}-4c_{0}+3\right)}{12k^{4}-7k^{2}-4c_{0}+3}AB,\quad A_{02}=-\frac{\left(4k^{2}c_{0}+4c_{0}+1\right)}{k^{2}-4c_{0}-1}A\overline{B}.$
For the Lagrange multiplier expansion we get
$\alpha_{2}=\frac{k^{2}\left(k^{2}-8c_{0}+6\right)}{2}{|A|}^{2}+\frac{k^{2}\left(k^{2}-8c_{0}+6\right)}{2}{|B|}^{2}-\beta_{2},$
and altogether this gives the solvability condition (67) for $A$ and $B$ at $\varepsilon^{3}$ with
$a=$ $k^{2},$
$b_{1}=$ $\frac{5}{4}k^{6}+5k^{4}c_{0}+\frac{3}{4}k^{4}+9k^{2}c_{0}-\frac{33}{4}k^{2}-6c_{0}-\frac{3}{2}$
$+\frac{3\left(2k^{4}+4k^{2}c_{0}-6k^{2}-4c_{0}-1\right)\left(-\frac{1}{2}k^{4}+k^{2}c_{0}-\frac{13}{4}k^{2}-2c_{0}-\frac{1}{2}\right)}{4k^{4}+7k^{2}+4c_{0}+1},$
$b_{2}=$ $3k^{6}+4k^{4}c_{0}-6k^{4}-16k^{2}c_{0}-3k^{2}-12c_{0}-3$
$+\frac{6\left(4k^{2}c_{0}+4c_{0}+1\right)\left(-\left(c_{0}+\frac{1}{4}\right)k^{2}+\frac{k^{2}}{2}-\frac{1}{2}-2c_{0}\right)}{k^{2}-4c_{0}-1}$
$+\frac{2\left(6k^{4}-4k^{2}c_{0}-4k^{2}-4c_{0}+3\right)k^{2}\left(-\frac{3k^{2}}{2}-c_{0}+\frac{3}{4}\right)}{12k^{4}-7k^{2}-4c_{0}+3}.$
Appendix B Numerical algorithms
@@
approximates the (positive definite) Laplace–Beltrami operator, while for the FEM mass matrix $M$ we use the so–called Voronoi mass matrix. Then, e.g., $MH(X)=\frac{1}{2}\left\langle LX,N\right\rangle$, where the vertex normals $N$ are obtained from interpolating the normals of the adjacent triangles. For the Gaussian curvature $K$ we use a discrete version of Gauss–Bonnet, and refer to [BGN19] for more sophisticated schemes. See [MU24b, §2.2] for comments on and illustrations of convergence of the used FEM.
Technically, for the 4th order problem (5) we use a mixed formulation as two 2nd order problems, first analyzed in [Rus05] (in a more complicated scheme); see [SGJW22] for a convergence analysis in the frame of the biharmonic equation. Additionally to the PDE $\eta(u){\partial}_{t}u=G(u,\Lambda)$ and the constraints $q(u)=0$, the continuation needs phase conditions (PCs), for which we introduce additional Lagrange multipliers $s_{x},s_{y},s_{z}$ and $s_{rx}$ pertaining to translations in $x,y$ and $z$, respectively, and rotations around the xx–axis. The yy–and zz–translation PCs are always on, the xx–translations is on for pearling and coiling, the xx–rotation is on for the wrinkling, and xx–translation and rotation are on for the buckling. In all continuations, the associated Lagrange multipliers stay ${\cal O}(10^{-6})$. For the discrete systems, an effect of all the constraints (also of $q(u)=0$) is that the linear systems (to be solved in Newton loops and for eigenvalue computations) are bordered, with border widths up to six rows/columns, and we usually use bordered elimination [Uec21, §4.5.1] for their solution.
Mesh handling
The initial cylinder in discretized by a standard mesh from pde2path, mapped via a local map on the cylinder. The number of nodes in $x$ and $\varphi$ direction is chosen to minimize the mesh–quality quotient
$q=\max_{T\in{\cal T}}\frac{e_{M}(e_{1}+e_{2}+e_{3})}{2|T|},$ (78)
with $e_{j}$ the edge lengths, $e_{M}$ the maximal edge length, and $|T|$ the triangle areas. The triangulation is oriented inwards, matching with our analytical convention. On the bifurcating branches triangles distort, due to the change of X. We use mesh adaptation by refinement and coarsening of X and tri based on (78) and triangle areas. Namely, we refine if the area is above a threshold, typically 1.3 times the size of the original cylinder triangles. For the conservation of the periodic boundary, we enforce the same refinement at “opposite” boundary triangles. For the coarsening (78) is the control parameter, activated if $q>8$. We select triangles smaller than a threshold $A_{t}$, typically $A_{t}=\max_{T\in{\cal T}}|T|/3$ and collapse acute triangles along the shortest edge, while obtuse triangles are handled with edge flipping.
Periodicity in $x$
During the continuation of bifurcating branches, it is not necessary that the boundary of $X$ stays in a plane. However, for mesh refinement along the branch the legacy box2per setting looks for periodic boundary nodes in planes. For that reason we only allow an $(y,z)$ displacement at the boundary, by setting $N_{x}=0$.
Extracting the amplitude
For the comparison between the numerics and the AE predictions, we evaluate the deviation at each mesh point. In each continuation step we obtain $X+uN$ as the new solution, but multiple steps result in tangential movement of mesh points relative to the cylinder. Therefore we project each mesh point back onto the cylinder orthogonally to $Cp=\left(p_{1},\frac{p_{2}}{\|(p_{2},p_{3})\|},\frac{p_{3}}{\|(p_{2},p_{3})\|}\right)^{T}.$ From this projection, we can measure the normal deviation from the cylinder and compare it with the amplitude approximation.
On the volume in the numerics
To compute the volume of (not necessarily straight) cylinders $\cal C$} we split them into two domains $\Omega_{1}$ and $\Omega_{2}$, see Fig.15, where $\Omega_{2}=\Omega_{2,a}\cup\Omega_{2,b}$ consists of two equal cones, from $(x,y,z)=(0,0,0)$ to the (in general not circular) left and right disk at $x=\pm L/2$. The idea is to compute $V={\rm vol}(\Omega_{1})+2{\rm vol}(\Omega_{2a})$ by the divergence theorem ${\rm vol}(\Omega)=\frac{1}{3}\int_{\Omega}\mathrm{div}(x)\,{\rm d}x=\frac{1}{3}\int_{{\partial}\Omega}\left\langle x,N\right\rangle\,{\rm d}S$. In the splitting ${\cal C}=\Omega_{1}\cup\Omega_{2}$, we have $\left\langle x,N\right\rangle=0$ along the inner “conical” boundaries ${\partial}{\Omega_{1}}\cap{\partial}{\Omega_{2}}$, such that ${\rm vol}(\Omega_{1})=\frac{1}{3}\int_{X}\left\langle(x,y,z),N\right\rangle\,{\rm d}S$, and ${\rm vol}(\Omega_{2a})=\frac{1}{3}\int_{X_{\text{left}}}\left\langle(-L/2,0,0)^{T},(-1,0,0)^{T}\right\rangle\,{\rm d}(y,z)=\frac{L}{2}|X_{\text{left}}|$, and for the latter we can apply Gauss’ theorem again to obtain
${\rm vol}(\Omega_{2a})=\frac{L}{6}\frac{1}{2}\int_{{\partial}X_{\text{left}}}\left\langle(y,z),(\nu_{1},\nu_{2})\right\rangle\,{\rm d}s=\frac{L}{12}\int_{{\partial}X_{\text{left}}}\left\langle(X_{2},X_{3}),(N_{2},N_{3})/\|(N_{2},N_{3})\|\right\rangle\,{\rm d}s.$ (79)
Thus, altogether,
${\rm vol}({\cal C})=\frac{1}{3}\int_{X}\left\langle(x,y,z),N\right\rangle\,{\rm d}S+\frac{L}{6}\int_{{\partial}X_{\text{left}}}\left\langle(X_{2},X_{3}),(N_{2},N_{3})/\|(N_{2},N_{3})\|\right\rangle\,{\rm d}s.$ (80)
Figure 15: Sketch for volume computation.
In the numerics we also need the derivative of $V({\cal C}+uN)$ wrt $u$. For $\Omega=\Omega_{0}+uN$ and ${\rm vol}(\Omega)=\frac{1}{3}\int_{{\partial}\Omega}\left\langle X,N\right\rangle\,{\rm d}S$ we have ${\partial}_{u}{\rm vol}(\Omega)v=\int_{\Omega}v\,{\rm d}S$, which can be used for ${\partial}_{u}{\rm vol}(\Omega_{1})$, and for ${\partial}_{u}{\rm vol}(\Omega_{2a})$ we proceed as follows. We choose a positive oriented arc length parametrization $\gamma:[0,2\pi)\rightarrow{\partial}X_{\text{left}}$. Then the (unit) tangent and (unit) normal vectors are $T(s)=\begin{pmatrix}\gamma_{1}^{\prime}(s),\gamma_{2}^{\prime}(s)\end{pmatrix}^{T}$ and $\nu(s)=\begin{pmatrix}-\gamma_{2}^{\prime}(s),\gamma_{1}^{\prime}(s)\end{pmatrix}^{T}$. For a perturbation $u$ we consider $\tilde{\gamma}(s)=\gamma(s)+\delta u(s)\nu(s)=\begin{pmatrix}\gamma_{1}(s)-\delta u(s)\gamma^{\prime}_{2}(s)\\ \gamma_{2}(s)+\delta u(s)\gamma^{\prime}_{1}(s)\end{pmatrix}$, with unit normal vector $\tilde{\nu}=\begin{pmatrix}-\tilde{\gamma}^{\prime}_{2}(s)\\ \tilde{\gamma}^{\prime}_{1}(s)\end{pmatrix}/\|\tilde{\gamma}^{\prime}(s)\|$. Then
$\frac{\,{\rm d}}{\,{\rm d}\delta}\bigg{|}_{\delta=0}$ $\int_{0}^{2\pi}\left\langle\tilde{\gamma}(s),\tilde{\nu}(s)\right\rangle\|\tilde{\gamma}^{\prime}(s)\|\,{\rm d}s=\frac{\,{\rm d}}{\,{\rm d}\delta}\bigg{|}_{\delta=0}\int_{0}^{2\pi}\tilde{\gamma}_{2}(s)\tilde{\gamma}^{\prime}_{1}(s)-\tilde{\gamma}_{1}(s)\tilde{\gamma}^{\prime}_{2}(s)\,{\rm d}s$
$=\int_{0}^{2\pi}\frac{\,{\rm d}}{\,{\rm d}\delta}\bigg{|}_{\delta=0}\Big{[}\big{(}\gamma_{2}(s)+\delta u(s)\gamma_{1}^{\prime}(s)\big{)}\Big{(}\gamma_{1}^{\prime}(s)-\delta\big{(}u(s)\gamma_{2}^{\prime}(s)\big{)}^{\prime}\Big{)}$
${\hskip 42.67912pt}-\big{(}\gamma_{1}(s)-\delta u(s)\gamma_{2}^{\prime}(s)\big{)}\Big{(}\gamma_{2}^{\prime}(s)+\delta\big{(}u(s)\gamma_{1}^{\prime}(s)\big{)}^{\prime}\Big{)}\Big{]}\,{\rm d}s$
$=\int_{0}^{2\pi}u(s)\gamma_{1}^{\prime}(s)^{2}-\gamma_{2}(s)(u(s)\gamma_{2}^{\prime}(s))^{\prime}+u(s)\gamma_{2}^{\prime}(s)^{2}-\gamma_{1}(s)(u(s)\gamma_{1}^{\prime}(s))^{\prime}\,{\rm d}s$
$=2\int_{0}^{2\pi}u(s)\,{\rm d}s,$
which can be easily implemented using the (1D) mass matrix of the discretization of ${\partial}X_{\text{left}}$.
Numerical flows
We give a brief description of our ad hoc method for numerical flows (aka direct numerical simulation DNS) for (5), and refer to [BMN05, BGN08, BGN20, BGN21] and the references therein for more sophisticated and provably stable FEM schemes for geometric flows, including the Helfrich flow (5) and extensions, and in particular also dealing with tangential motion of mesh points. Here we only consider normal displacements, and approximate the normal velocity as
$V_{t}=\left\langle{\partial}_{t}X,N\right\rangle\approx\left\langle\frac{1}{h}M_{n}(X_{n+1}{-}X_{n}),N_{n}\right\rangle=\frac{1}{h}\left\langle M_{n}(X_{n}{+}u_{n+1}N_{n}{-}X_{n}),N_{n}\right\rangle=\frac{1}{h}M_{n}u_{n+1},$
where $h$ is the current stepsize, $n$ the time stepping index, and $M_{n}$ the mass matrix. Setting $U_{n}=(u_{n},\Lambda_{n})$, the PDE $V_{t}=G(X(t),\Lambda(t))$, together with the (volume and area) constraints $q(u_{n+1})=0$ (translational and rotational phase conditions play no role in the DNS), yields the nonlinear system
${\cal G}(U_{n+1}):=\begin{pmatrix}\frac{1}{h}M_{n}u_{n+1}+G(U_{n+1})\\ q(u_{n+1})\end{pmatrix}\stackrel{{\scriptstyle!}}{{=}}\begin{pmatrix}0\\ 0\end{pmatrix}\in{\cal H}\times{\mathbb{R}}^{n_{q}}$ (81)
@@
Additionally we remark that Newton loops for (81) often fail to converge near straight cylinders (or more generally CMC surfaces), where ${\partial}_{u}q_{1}$ and ${\partial}_{u}q_{2}$ become linearly dependent ($\cal D$} from (20) becomes singular), but this is not a major concern for us as we are more interested in the flow towards nontrivial steady states. See, e.g., [BGN08, Scheme (2.8)] for dealing with the constraints when approaching CMC surfaces, which essentially consists in setting $\lambda_{1,n+1}=\lambda_{1,n}$ and only solving for $\lambda_{2,n+1}$ with $q(u_{n+1})=q_{2}(u_{n+1})$. Moreover, the DNS in §5 are all uncritical wrt to mesh quality. However, for stronger initial perturbations we expect that allowing and controlling tangential motion of mesh points (or other ways of mesh–adaptation) will be crucial, see also, e.g., the numerical flows for closed vesicles in [MU24a].